8 January 2009 Lab Meeting Agenda
- Check with B&G about compressed air and vacuum?
- Meeting Time for Spring Semester is THURSDAYS 3-5pm ROOM 705. Stating 1/8/09. See Lab Calendar for dates and presentation schedule.
† Flu R01
- Check on sample amount.
- Plaque Assay.
- Alex needs to be trained to make SU molds. Trained to make the monoliths (by Hussam By 1/15/09)--we're about to start the training--.
† SEPSIS Project
- Sonali working on first draft of new version paper by 1/19.
- Hussam working on Lambda phage data.
- 1 ml methanol wash seems to be taking care of the issue.
- More formal report to Jeff from Cathie.
- Do cDNA prep. Spec, gels, send to Jeff B. Shipping?
- Mark- Control circuit - This will be done by January.
- Evap quantification experiments ongoing. Still working on live/dead assay.(JD, JZ)
- problems: dead but green in the end
- called tech support.
- tests: supernatant fluorescent? excessive dye to observe time-lapse change in green/red ratio.
- omit washing step before dilution: wash, dye, incubate, dilute, instead of wash, dye, incubate, wash, dilute
- try running the experiment first, then running the live/dead dye through second. Bold text
- Petridish culture after running evaporation w/o dye in hydrophobic treated channel.
- Hydrophobic coating of microchannel
- RIE (CHF3), Rain?, and Cytop coated and thermal bonding with PTFE filter tested.
- Cytop turned out good bonding with PTFE filter.
- Test cheap wait for running
- Slide for dispensing or integration scheme with SERS substrate
- new set made, in glove box: the 1st 3 on the DOE. SEM and SERS available.
- hypothesis: drying controls clustering, reduction controls size?
- need to add humidity as it changes even in the glove box - this is no longer an issue, asked Ranjith and now can adjust humidity to constant
- set II done varying drying time and fast reduction time. SEM and SERS available.
- Colloid experiment. inconclusive. Look at SEMs
- difficulty in locating the bacteria
- needs much higher particle/bacteria ratio
- experiments outside the channel first?
- Look at the unused substrates in SEM one week old or more.
- Prospectus framework/discussion: target the end of this week 1/23/09.
† Valve Array/Valve building (FJ)
† Fraunhofer: LOAC
† Biointerfaces group
- Strobe experiments? Jason cannot do them, perhaps MCK can. MCK will talk to Jason this week. Look at the bacteria as they flow.
- CR mask was printed new design for trapper .
- UV filter installed
- Cathie needs a copy of MCK's, Journal of Applied Physics paper.
- Experimental Data for paper.What do we need in terms of flow rate, seeding densities, etc.
- Do 10 micron simulations.
† CIMIT- Colson Grant
- IRB still going/ revisions by tomorrow to Ian. Target date 1/1/09
- STAR-CD Renewal - January ($2K)
- Only 1/5 chips are good at the hot embossing step. The vaccum chuck will be checked and o-rings changed - helped a bit. ON HOLD
* Fixuture for PCR chip. The heater and the thermocouples would be on the fixture. ON HOLD
- Try microscope for looking a fluor. Learn how to look at the data in Image J
- Cathie submitted One pager Whitepaper. Waiting for response.
† Senior project
- Megan 1)did some preliminary bonding with Zeonex and teflon film. 2) emailed off the Mask Design to FineLine
† F31: Cochlea
- Set up adviser meeting Jan
- Post Gent vs. Pre gent - 6 proteins. 4/6 no change.
- Final two proteins: know by week of 1/19
† Silica Optimization (Lambda):
- Elution of C and D total DNA mass 2ng and 0.2 ng 
- Testing with water 12 channels- standard protocol- NA replaced with NF water 
- 10 channels - 5ch methanol-water / 5ch methanol-buffer-water 
- Next E. coli.